OpenStudy (anonymous):

given that vector OA = 4i+3j and C is a point on OA such that the magnitude of vector OC is 16/5. a/ find vector OC. b/ is vector OB = i+4j, show that BC is perpendicular to OA.

OpenStudy (anonymous):

a) since C is on OA so it's directiion is same as OA so only magnitude is changing so OC=(OA/|OA|)*(16/5)=(64/25)i+(48/25)j b) given B(1,4) & C((64/25),(48/25)) so BC=((64/25)-1)i+((48/25)-4)j=(39/25)i-(52/25)j now if BC.OA=0 then we can say that OA and BC are perpendicular BC.OA=(156/25)-(156/25)=0 hence proved